OpenStudy (anonymous):

f(x) = 5x + 3; g(x) = 6x - 5 Find f/g. and the domain

OpenStudy (anonymous):

\[\frac{ f }{ g }=\frac{ 5x+3 }{ 6x-5 },domain is 6x-5\neq0, x \neq \frac{ 5 }{ 6 }\] domain is all reals except 5/6